OpenStudy (anonymous):

okay, I feel so stupid for forgetting how to do this...But what happens next? It's; 4x+5y=20. I know I subtract 5 y on each sides, but what happens next?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (shamil98):

Are you converting to slope-intercept form?

OpenStudy (shamil98):

If so, 4x + 5y = 20 subtract both sides by 4x then divide both sides by 5 to get slope-intercercept form y = mx + b

OpenStudy (anonymous):

Oh, yeah.. I think the answer is y=.8x+4? Do I do something next?

OpenStudy (anonymous):

@shamil98

OpenStudy (shamil98):

5y = -4/5 x + 4 you leave -4/5 (the slope) as it is -4/5 = down 4 and right 5 , and up 4 left 5. you start graphing it from (0,4) 4 is the y-intercept(b)